OpenStudy (firejay5):

write an equation in slope- intercept form for the line that satisfies each set of conditions. 30. passes through (7,1) and (7,8) 32. passes through(-2, -3) and (0,0)

OpenStudy (anonymous):

y_2-y_1=m(x_2-X_1)

OpenStudy (firejay5):

I need the answer in y intercept for

OpenStudy (campbell_st):

can you find the slope of the line..?

OpenStudy (firejay5):

I do, but don't have time to do it and the teacher wants answers only no work shown

OpenStudy (campbell_st):

ok... if you know the slope use y = mx + b substitute the x and y values from a point to evaluate b

OpenStudy (firejay5):

at least I will give you a medal for it

OpenStudy (campbell_st):

well if you don't have time to find the slope you can't find the equation

OpenStudy (firejay5):

if you tell me the steps

OpenStudy (firejay5):

can you just tell me the answer

OpenStudy (campbell_st):

whats the slope in 30...? 32 has a slope of 3/2

OpenStudy (firejay5):

30 is undefined in slope

OpenStudy (campbell_st):

for 32. y = 3/2 x + b use the point (0, 0) x = 0 and y = 0 to find b, substitute 0 = 3/2* 0 + b b = 0 therefore the equation is y = 3/2x just do the same for 30, you said you knew the slope

OpenStudy (campbell_st):

so what which lines have an undefined slope..?

OpenStudy (firejay5):

x

OpenStudy (firejay5):

what's the y - intercept